If Jane volunteered at the hospital 12 hours in January and each day she works 4/5 an hour how many days does she volunteer in January
Answer by stanbon(75887)

If Jane volunteered at the hospital 12 hours in January and each day she works 4/5 an hour how many days does she volunteer in January.
---------------------------
Ans: 12/(4/5) = 3*5 = 15 days
--------------
Cheers,
Stan H.
------------------